# A Survey Analyzing Generalization in Deep Reinforcement Learning

### 04 Jan 2024 v1

@@ 6.3 The Adversarial Perspective for Deep Neural Policy Generalization
One of the ways to regularize the state observations is based on considering worst-case perturbations added to state observations (i.e. adversarial perturbations). This line of work starts with introducing perturbations produced by the fast gradient sign method proposed by Goodfellow et al. (2015) into deep reinforcement learning observations at test time Huang et al. (2017) Kos & Song (2017), and compares the generalization capabilities of the trained deep reinforcement learning policies in the presence worst-case perturbations and Gaussian noise. These gradient based adversarial methods are based on taking the gradient of the cost function used to train the policy with respect to the state observation. Several other techniques have been proposed on the optimization line of the adversarial alteration of state observations. In this line of work, Korkmaz (2020) suggested a Nesterov momentum-based method to produce adversarial perturbations for deep reinforcement learning policies. Korkmaz (2022) further showed that deep reinforcement learning policies learn shared adversarial features across MDPs. In this work the authors investigate the root causes of this problem, and demonstrate that policy high-sensitivity directions and the perceptual similarity of the state observations are uncorrelated. Furthermore, the study demonstrates that the current state-of-the-art adversarial training techniques also learn similar high-sensitivity directions as the vanilla trained deep reinforcement learning policies. [^4] From the security point of view, this adversarial framework is under the category of black-box adversarial attacks for which this is the first study that demonstrated that deep reinforcement learning policies are vulnerable to black-box adversarial attacks (Korkmaz, 2022). Furthermore, note that black-box adversarial perturbations are more generalizable global perturbations that can effect many different policies. While some studies focused on state observation alterations to assess policy resilience with respect to these changes, some studies focused on interpretability and explainability of these changes in these state observation alterations and how these alterations have different effects on standard deep reinforcement learning training algorithms and certified robust (i.e. adversarial) training algorithms Korkmaz (2021e) [^5] See an initial and preliminary version of this paper here Korkmaz (2021a; d).
Table 4: Environment and algorithm details for adversarial policy regularization and attack techniques in deep reinforcement learning. Note that most of the methods based on adversarial policy regularization are a form of state observation perturbation method $\mathcal{G}_{S}^{S}$ as described in Definition 3.6.
Citation Proposed Method Environment Reinforcement Learning Algorithm
Huang et al. (2017) FGSM ALE DQN, TRPO, A3C
Kos & Song (2017) FGSM ALE DQN and IQN
Lin et al. (2017) Strategically-Timed Attack ALE A3C and DQN
Gleave et al. (2020) Adversarial Policies Mujoco Proximal Policy Optimization
Huan et al. (2020) SA-DQN ALE and $L_{\textrm{Mujoco}}$ [^6] Low dimensional state Mujoco refers to the setting of Mujoco where the state dimensions are not represented by pixels and dimensions of the state observations range from 11 to 117. DDQN and PPO
Korkmaz (2022) Adversarial Framework ALE DDQN and A3C
Korkmaz (2023) Natural Attacks ALE DDQN and A3C
Korkmaz & Brown-Cohen (2023) Adversarial Detection ALE DDQN
Note that this line of work falls under the state observation generalization testing category $\mathcal{G}_{S}^{S}$ provided in Definition 3.6.
@@
While several studies focused on improving optimization techniques to compute optimal perturbations, a line of research focused on making deep neural policies resilient to these perturbations. Pinto et al. (2017) proposed to model the dynamics between the adversary and the deep neural policy as a zero-sum game where the goal of the adversary is to minimize expected cumulative rewards of the deep reinforcement learning policy. This study is a clear example of transition probability perturbation to achieve generalization $\mathcal{G}_{\mathcal{P}}$ in Definition 3.4 of Section 3. Gleave et al. (2020) approached this problem with an adversary model which is restricted to take natural actions in the MDP instead of modifying the observations with $\ell_{p}$-norm bounded perturbations. The authors model this dynamic as a zero-sum Markov game and solve it via self play Proximal Policy Optimization (PPO). Some recent studies, proposed to model the interaction between the adversary and the deep reinforcement learning policy as a state-adversarial MDP, and claimed that their proposed algorithm State Adversarial Double Deep Q-Network (SA-DDQN) learns theoretically certified robust policies against natural noise and perturbations. In particular, these certified adversarial training techniques aim to add a regularizer term to the temporal difference loss in deep $Q$-learning.
@@
where $\mathcal{H}$ is the Huber loss, $\hat{Q}$ refers to the target network and $\kappa$ is to adjust the level of regularization for convergence. The regularizer term can vary for different certified adversarial training techniques yet the baseline technique uses
$\mathcal{R}(\theta)=\max\{\max_{\hat{s}\in B(s)}\max_{a\neq\operatorname*{arg\,max}_{a^{\prime}}Q(s,a^{\prime})}Q_{\theta}(\hat{s},a)-Q_{\theta}(\hat{s},\operatorname*{arg\,max}_{a^{\prime}}Q(s,a^{\prime}),-c\}.$
where $B(s)$ is an $\ell_{p}$-norm ball of radius $\epsilon$. While these certified adversarial training techniques drew some attention from the community, more recently manifold concerns have been raised on the robustness of theoretically certified adversarially trained deep reinforcement learning policies (Korkmaz, 2021e; 2022). In these studies, the authors argue that adversarially trained (i.e. certified robust) deep reinforcement learning policies learn inaccurate state-action value functions and non-robust features from the environment. In particular, in Korkmaz (2021c) the authors use action manipulation to investigate worst-case perturbation training. This study is also a clear example of a policy perturbation generalization testing method $\mathcal{G}_{S}^{\pi}$ in Definition 3.6. More importantly, recently it has been shown that adversarially trained deep reinforcement learning policies have worse generalization capabilities compared to vanilla trained reinforcement learning policies in high dimensional state space MDPs (Korkmaz, 2023) [^7] A short and preliminary version of the paper (Korkmaz, 2023) can also be found here (Korkmaz, 2021b). While this study provides a contradistinction between adversarial directions and natural directions that are intrinsic to the MDP, it further demonstrates that the certified adversarial training techniques block generalization capabilities of standard deep reinforcement learning policies. Furthermore note that this study is also a clear example of a state observation perturbation generalization testing method $\mathcal{G}_{S}^{S}$ in Definition 3.6 in Section 3.
@@ 7 Meta-Reinforcement Learning and Meta Gradients
A quite recent line of research directs its research efforts to discovering reinforcement learning algorithms automatically, without explicitly designing them, via meta-gradients (Oh et al., 2020; Xu et al., 2020). This line of study targets learning the ”learning algorithm” by only interacting with a set of environments as a meta-learning problem. In particular,
@@
here the optimal update rule is parametrized by $\eta$, for a distribution on environments $\rho(\varepsilon)$ and initial policy parameters $\rho(\theta_{0})$ where $\mathbb{E}_{\theta_{N}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ is the expected return for the end of the lifetime of the agent.
The objective of meta-reinforcement learning is to be able to build agents that can learn how to learn over time, thus allowing these policies to adapt to a changing environment or even any other changing conditions of the MDP. Quite recently, a significant line of research has been conducted to achieve this objective, particularly Oh et al. (2020) proposes to discover update rules for reinforcement learning. This line of work also falls under the policy transformation generalization $\mathcal{G}_{\pi}$ in Definition 3.5 defined in Section 3. Following this work Xu et al. (2020) proposed a joint meta-learning framework to learn what the policy should predict and how these predictions should be used in updating the policy. Recently, Kirsch et al. (2022) proposes to use symmetry information in discovering reinforcement learning algorithms and discusses meta-generalization.
There is also some work on enabling reinforcement learning algorithms to discover temporal abstractions (Veeriah et al., 2021). In particular, temporal abstraction refers to the ability of the policy to abstract a sequence of actions to achieve certain sub-tasks. As it is promised within this subfield, meta-reinforcement learning is considered to be a research direction that could enable us to build deep reinforcement learning policies that can generalize to different environments, to changing environments over time, or even to different tasks.
@@ 8 Transfer in Reinforcement Learning
Transfer in reinforcement learning is a subfield heavily discussed in certain applications of reinforcement learning algorithms e.g. robotics. In current robotics research there is not a safe way of training a reinforcement learning agent by letting the robot explore in real life. Hence, the way to overcome this is to train policies in a simulated environment, and install the trained policies in the actual application setting. The fact that the simulation environment and the installation environment are not identical is one of the main problems for reinforcement learning application research. This is referred to as the sim-to-real gap.
Another subfield in reinforcement learning research focusing on obtaining generalizable policies investigates this concept through transfer in reinforcement learning. The consideration in this line of research is to build policies that are trained for a particular task with limited data and to try to make these policies perform well on slightly different tasks. An initial discussion on this starts with (Taylor & Stone, 2007) to obtain policies initially trained in a source task and transferred to a target task in a more sample efficient way. Later, Tirinzoni et al. (2018) proposes to transfer value functions that are based on learning a prior distribution over optimal value functions from a source task. However, this study is conducted in simple environments with low dimensional state spaces. (Barreto et al., 2017) considers the reward transformation setting $\mathcal{G}_{R}$ in Definition 3.2 from Section 3. In particular, the authors consider a policy transfer between a specific task with a reward function $r(s,a)$ and a different task with reward function $r^{\prime}(s,a)$. The goal of the study is to decouple the state representations from the task. In the setting of state transformation for generalization $\mathcal{G}_{S}$ in Definition 3.3 Gamrian & Goldberg (2019) focuses on state-wise differences between source and target task. In particular, the authors use unaligned generative adversarial networks to create target task states from source task states. In the setting of policy transformation for generalization $\mathcal{G}_{\pi}$ in Definition 3.5 Jain et al. (2020) focuses on zero-shot generalization to a newly introduced action set to increase adaptability.
While transfer learning is a promising research direction for reinforcement learning, the studies in this subfield still remain oriented only towards reinforcement learning applications, and it is possible to consider the research centered on this subfield as not at the same level of maturity as the previously discussed line of research in Section 6 in terms of being able to test the claims or propositions in complex established baselines.
@@ 9 Lifelong Reinforcement Learning
Lifelong learning is a subfield closely related to transfer learning that has recently drawn attention from the reinforcement learning community. Lifelong learning aims to build policies that can sequentially solve different tasks by being able to transfer knowledge between tasks. On this line of research, Lecarpentier et al. (2021) provide an algorithm for value-based transfer in the Lipschitz continuous task space with theoretical contributions for lifelong learning goals. In the setting of action transformation for generalization $\mathcal{G}_{\pi}$ in Definition 3.5 Chandak et al. (2020) focuses on temporally varying (e.g. variations between source task and target task) the action set in lifelong learning. In lifelong reinforcement learning some studies focus on different exploration strategies. In particular, Garcia & Thomas (2019) models the exploration strategy problem for lifelong learning as another MDP, and the study uses a separate reinforcement learning agent to find an optimal exploration method for the initial lifelong learning agent. The lack of benchmarks limits the progress of lifelong reinforcement learning research by restricting the direct comparison between proposed algorithms or methods. However, quite recent work proposed a new training environment benchmark based on robotics applications for lifelong learning to overcome this issue (Wolczyk et al., 2021) [^8] The state dimension for this benchmark is 12. Hence, the state space is low dimensional..
@@ 10 Inverse Reinforcement Learning
Inverse reinforcement learning focuses on learning a functioning policy in the absence of a reward function. Since the real reward function is inaccessible in this setting and the reward function needs to be learnt from observing an expert completing the given task, the inverse reinforcement learning setting falls under the reward transformation for generalization setting $\mathcal{G}_{R}$ defined in Definition 3.2 in Section 3. The initial work that introduced inverse reinforcement learning was proposed by Ng & Russell (2000) demonstrating that multiple different reward functions can be constructed for an observed optimal policy. The authors of this initial study achieve this objective via linear programming,
$\textrm{maximize}\sum_{s\in S_{\rho}}$ $\min_{a\in A}\{p(\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,a_{1}|\cdot)}\mathcal{V}^{\pi}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,a|\cdot)}\mathcal{V}^{\pi}(s^{\prime}))\}$ (2)
@@
where $p(x)=x$ if $x\geq 0$, $p(x)=2x$ otherwise and $\mathcal{V}^{\pi}=\alpha_{1}\mathcal{V}^{\pi}_{1}+\alpha_{2}\mathcal{V}^{\pi}_{2}+\dots+\alpha_{d}\mathcal{V}^{\pi}_{d}$. In this line of work, there has been recent progress that achieved learning functioning policies in high-dimensional state observation MDPs (Garg et al., 2021). The study achieves this by learning a soft $Q$-function from observing expert demonstrations, and the study further argues that it is possible to recover rewards from the learnt soft state-action value function.

### 30 Oct 2024 v2

@@ 6.3 The Adversarial Perspective for Deep Neural Policy Generalization
One of the ways to regularize the state observations is based on considering worst-case perturbations added to state observations (i.e. adversarial perturbations). This line of work starts with introducing perturbations produced by the fast gradient sign method proposed by Goodfellow et al., (2015) into deep reinforcement learning observations at test time (Huang et al.,, 2017; Kos and Song,, 2017), and compares the generalization capabilities of the trained deep reinforcement learning policies in the presence worst-case perturbations and Gaussian noise. These gradient based adversarial methods are based on taking the gradient of the cost function used to train the policy with respect to the state observation.
$\mathnormal{\displaystyle s_{\textrm{adv}}=s+\epsilon\cdot\frac{\nabla_{x}J(\displaystyle s,Q(s,a))}{\left\lVert\nabla_{s}J(s,Q(s,a))\right\rVert_{p}},}$
Several other techniques have been proposed on the optimization line of the adversarial alteration of state observations. In this line of work, Korkmaz, (2020) suggested a Nesterov momentum-based method to produce adversarial perturbations for deep reinforcement learning policies.
$v_{t+1}=\mu\cdot v_{t}$ $+\dfrac{\nabla_{s_{\textrm{adv}}}J(s_{\textrm{adv}}^{t}+\mu\cdot v_{t},a)}{\lVert\nabla_{s_{\textrm{adv}}}J(s_{\textrm{adv}}^{t}+\mu\cdot v_{t},a)\rVert_{1}}$
$s_{\textrm{adv}}^{t+1}$ $=s_{\textrm{adv}}^{t}+\alpha\cdot\dfrac{v_{t+1}}{\lVert v_{t+1}\rVert_{2}}$
Here $J(s_{\textrm{adv}},a)$ is based on the cost function used to train the policy, $s_{\textrm{adv}}$ represents the adversarial state observation, and $\mu$ is the momentum acceleration parameter. While a line of studies focused on optimization aspects of the adversarial perturbations, some studies demonstrated further the hidden linearity of deep reinforcement learning policies by revealing how these policies learn shared adversarial features across states, MDPs and across algorithms (Korkmaz,, 2022).
@@
In this work the authors investigate the root causes of this problem, and demonstrate that policy high-sensitivity directions and the perceptual similarity of the state observations are uncorrelated. Furthermore, the study demonstrates that the current state-of-the-art adversarial training techniques also learn similar high-sensitivity directions as the vanilla trained deep reinforcement learning policies. [^5] From the security point of view, this adversarial framework is under the category of black-box adversarial attacks for which this is the first study that demonstrated that deep reinforcement learning policies are vulnerable to black-box adversarial attacks (Korkmaz,, 2022). Furthermore, note that black-box adversarial perturbations are more generalizable global perturbations that can affect many different policies. More recently, a line of work proposed theoretically founded algorithms to understand the temporal and spatial correlation of deep reinforcement learning decision making and what affects this decision making process (Korkmaz, 2024c, ). In particular, this study identifies what precisely affects and contributes to the decision making process of deep reinforcement learning policies from distributional shift to worst-case perturbations (i.e. adversarial), from algorithmic differences to architectural changes.
While several studies focused on improving computation techniques to optimize optimal perturbations, a line of research focused on making deep neural policies resilient to these perturbations. Pinto et al., (2017) proposed to model the dynamics between the adversary and the deep neural policy as a zero-sum game (Littman,, 1994) where the goal of the adversary is to minimize expected cumulative rewards of the deep reinforcement learning policy.
$R^{\textrm{agent}}=\mathbb{E}_{s_{0}\sim\rho;a^{\textrm{agent}}\sim\pi^{\textrm{agent}};a^{\textrm{adv}}\sim\pi^{\textrm{adv}}}[\sum_{t=0}^{T-1}r^{\textrm{agent}}(s,a^{\textrm{agent}},a^{\textrm{adv}})]$
[…]
[0pt]Refer to caption
Figure 1: Robust adversarial reinforcement learning proposed in (Pinto et al.,, 2017). This paper proposes the zero-sum game to model the relationship between the agent and the adversary while focusing on introducing disturbances to the environment dynamics. Here the empirical studies are conducted in the MuJoCo environment.
Here the adversarial policy is represented by $\pi^{\textrm{adv}}$, the policy of the agent represented by $\pi^{\textrm{agent}}$, and the rewards received by the agent represented by $r^{\textrm{agent}}$. The Nash equilibrium of the optimal rewards for this zero-sum game is
$R^{\textrm{agent}^{*}}=\min_{\pi^{\textrm{adv}}}\max_{\pi^{\textrm{agent}}}R^{\textrm{agent}}(\pi^{\textrm{agent}},\pi^{\textrm{adv}})=\max_{\pi^{\textrm{agent}}}\min_{\pi^{\textrm{adv}}}R^{\textrm{agent}}(\pi^{\textrm{agent}},\pi^{\textrm{adv}})$
@@
This study is a clear example of transition probability perturbation to achieve generalization $\mathcal{G}_{\mathcal{P}}$ in Definition 3.6 of Section 3. Gleave et al., (2020) approached this problem with an adversary model which is restricted to take natural actions in the MDP instead of modifying the observations with $\ell_{p}$-norm bounded perturbations. The authors model this dynamic as a zero-sum Markov game and solve it via self play Proximal Policy Optimization (PPO). Some recent studies, proposed to model the interaction between the adversary and the deep reinforcement learning policy as a state-adversarial MDP, and claimed that their proposed algorithm State Adversarial Double Deep Q-Network (SA-DDQN) learns theoretically certified robust policies against natural noise and perturbations. In particular, these certified adversarial training techniques aim to add a regularizer term to the temporal difference loss in deep $Q$-learning
@@
where $\mathcal{H}$ is the Huber loss, $\hat{Q}$ refers to the target network and $\kappa$ is to adjust the level of regularization for convergence. The regularizer term can vary for different certified adversarial training techniques yet the baseline technique uses $\mathcal{R}(\theta)$
$\max\{\max_{\hat{s}\in B(s)}\max_{a\neq\operatorname*{arg\,max}_{a^{\prime}}Q(s,a^{\prime})}Q_{\theta}(\hat{s},a)-Q_{\theta}(\hat{s},\operatorname*{arg\,max}_{a^{\prime}}Q(s,a^{\prime})),-c\}.$
where $B(s)$ is an $\ell_{p}$-norm ball of radius $\epsilon$. While these certified adversarial training techniques drew some attention from the community, more recently manifold concerns have been raised on the robustness of theoretically certified adversarially trained deep reinforcement learning policies (Korkmaz, 2021c, ; Korkmaz, 2021b, ; Korkmaz,, 2022; Korkmaz, 2024a, ). In these studies, the authors argue that adversarially trained (i.e. certified robust) deep reinforcement learning policies learn inaccurate state-action value functions and non-robust features from the environment. More importantly, recently it has been shown that certified robust deep reinforcement learning policies have worse generalization capabilities compared to vanilla trained reinforcement learning policies in high dimensional state space MDPs (Korkmaz,, 2023). While this study provides a contradistinction between adversarial and natural directions that are intrinsic to the MDP, it further demonstrates that the certified adversarial training techniques block generalization capabilities of standard deep reinforcement learning policies. Furthermore note that this study is also a clear example of a state observation perturbation generalization testing method $\mathcal{G}_{S}^{S}$ in Definition 3.8 in Section 3. For a more comprehensive view on generalization and robustness see Korkmaz, 2024b .
It is important to observe that the methods that focuses on improving generalization, i.e. robust training, described in this section rarely employ the different generalization testing methods proposed by other work. Thus, focusing narrowly on one aspect of generalization with one dimensional improvements in actuality decreases generalization on another aspect, as has been shown in the case of adversarial training (Korkmaz,, 2023). Therefore we again emphasize the need to understand the significance of a concrete definition of generalization, and a unified baseline to precisely measure it.
[…]
Figure 2: State transformation generalization under adversarial perspective in the Arcade Learning Environment (Korkmaz,, 2023). Note that under the adversarial influence direction of research, the state transformation generalization is constrained by the imperceptibility of the transformations. Columns: base frame, shifting, perspective transformation, blurring, discrete cosine transform artifacts, brightness and contrast. Up: JamesBond. Down: BankHeist.
@@ 7 Meta-Reinforcement Learning and Meta Gradients
A quite recent line of research directs its research efforts to discovering reinforcement learning algorithms automatically, without explicitly designing them, via meta-gradients (Oh et al.,, 2020; Xu et al.,, 2020). This line of study targets learning the ”learning algorithm” by only interacting with a set of environments as a meta-learning problem. In particular,
@@
here the optimal update rule is parametrized by $\eta$, for a distribution on environments $\rho(\varepsilon)$ and initial policy parameters $\rho(\theta_{0})$ where $\mathbb{E}_{\theta_{N}}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$ is the expected return for the end of the lifetime of the agent. The objective of meta-reinforcement learning is to be able to build agents that can learn how to learn over time, thus allowing these policies to adapt to a changing environment or even any other changing conditions of the MDP.
Quite recently, a significant line of research has been conducted to achieve this objective, particularly Oh et al., (2020) proposes to discover update rules for reinforcement learning. This line of work also falls under the algorithmic generalization $\mathcal{G}_{\mathbb{A}}$ in Definition 3.3 defined in Section 3. Following this work Xu et al., (2020) proposed a joint meta-learning framework to learn what the policy should predict and how these predictions should be used in updating the policy. Recently, Kirsch et al., (2022) proposes to use symmetry information in discovering reinforcement learning algorithms and discusses meta-generalization. There is also some work on enabling reinforcement learning algorithms to discover temporal abstractions (Veeriah et al.,, 2021). In particular, temporal abstraction refers to the ability of the policy to abstract a sequence of actions to achieve certain sub-tasks. As it is promised within this subfield, meta-reinforcement learning is considered to be a research direction that could enable us to build deep reinforcement learning policies that can generalize to different environments, to changing environments over time, or even to different tasks.
[…]
[0pt]Refer to caption
Figure 3: Meta training of the learned policy gradient that have been described in (Oh et al.,, 2020). Right: The learned policy gradient algorithm that has been trained in toy examples can generalize to more complex environment such as the Arcade Learning Environment.
@@ 8 Transfer in Reinforcement Learning
Transfer in reinforcement learning is a subfield heavily discussed in certain applications of reinforcement learning algorithms, e.g. robotics. In current robotics research there is not a safe way of training a reinforcement learning agent by letting the robot explore in real life. Hence, the way to overcome this is to train policies in a simulated environment, and install the trained policies in the actual application setting. The fact that the simulation environment and the installation environment are not identical is one of the main problems for reinforcement learning application research. This is referred to as the sim-to-real gap.
Another subfield in reinforcement learning research focusing on obtaining generalizable policies investigates this concept through transfer in reinforcement learning. The consideration in this line of research is to build policies that are trained for a particular task with limited data and to try to make these policies perform well on slightly different tasks. An initial discussion on this starts with Taylor and Stone, (2007) to obtain policies initially trained in a source task and transferred to a target task in a more sample efficient way. Later, Tirinzoni et al., (2018) proposes to transfer value functions that are based on learning a prior distribution over optimal value functions from a source task. However, this study is conducted in simple environments with low dimensional state spaces. Barreto et al., (2017) considers the reward transformation setting $\mathcal{G}_{R}$ in Definition 3.4 from Section 3. In particular, the authors consider a policy transfer between a specific task with a reward function $r(s,a)$ and a different task with reward function $r^{\prime}(s,a)$. The goal of the study is to decouple the state representations from the task. In the setting of state transformation for generalization $\mathcal{G}_{S}$ in Definition 3.5 Gamrian and Goldberg, (2019) focuses on state-wise differences between source and target task. In particular, the authors use unaligned generative adversarial networks to create target task states from source task states. In the setting of policy transformation for generalization $\mathcal{G}_{\pi}$ in Definition 3.7 Jain et al., (2020) focuses on zero-shot generalization to a newly introduced action set to increase adaptability. While transfer learning is a promising research direction for reinforcement learning, the studies in this subfield still remain oriented only towards reinforcement learning applications, and thus the main focus on applications centered on this subfield provides a non-unified progress in research due to the lack of an established baseline in which the proposed claims and algorithms can be consistently compared.
[…]
[0pt]Refer to caption
Figure 4: Transfer in reinforcement learning as has been described in (Gamrian and Goldberg,, 2019) that falls under the generalization through observation category explained in Definition 3.5. The frames are taken from Breakout game in the Arcade Learning Environment. The left frames represent the target task and the right frames represents the source tasks generated via generative adversarial networks.
@@ 9 Lifelong Reinforcement Learning
Lifelong learning is a subfield closely related to transfer learning that has recently drawn attention from the reinforcement learning community. Lifelong learning aims to build policies that can sequentially solve different tasks by being able to transfer knowledge between tasks. On this line of research, Lecarpentier et al., (2021) provide an algorithm for value-based transfer in the Lipschitz continuous task space with theoretical contributions for lifelong learning goals. In the setting of action transformation for generalization $\mathcal{G}_{\pi}$ in Definition 3.7 Chandak et al., (2020) focuses on temporally varying (e.g. variations between source task and target task) the action set in lifelong learning. In lifelong reinforcement learning some studies focus on different exploration strategies. In particular, Garcia and Thomas, (2019) models the exploration strategy problem for lifelong learning as another MDP, and the study uses a separate reinforcement learning agent to find an optimal exploration method for the initial lifelong learning agent. The lack of benchmarks limits the progress of lifelong reinforcement learning research by restricting the direct comparison between proposed algorithms or methods. However, quite recent work proposed a new training environment benchmark based on robotics applications for lifelong learning to overcome this issue (Wolczyk et al.,, 2021) [^6] The state dimension for this benchmark is 12. Hence, the state space is low dimensional..
@@ 10 Inverse Reinforcement Learning
Inverse reinforcement learning focuses on learning a functioning policy in the absence of a reward function. Since the real reward function is inaccessible in this setting and the reward function needs to be learnt from observing an expert completing the given task, the inverse reinforcement learning setting falls under the reward transformation for generalization setting $\mathcal{G}_{R}$ defined in Definition 3.4 in Section 3. The initial work that introduced inverse reinforcement learning was proposed by Ng and Russell, (2000) demonstrating that multiple different reward functions can be constructed for an observed optimal policy. The authors of this initial study achieve this objective via linear programming,
$\textrm{max}\sum_{s\in S_{\rho}}$ $\min_{a\in A}\{p(\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,a_{1}|\cdot)}\mathcal{V}^{\pi}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim\mathcal{P}(s,a|\cdot)}\mathcal{V}^{\pi}(s^{\prime}))\}$
@@
where $p(x)=x$ if $x\geq 0$, $p(x)=2x$ otherwise and $\mathcal{V}^{\pi}=\alpha_{1}\mathcal{V}^{\pi}_{1}+\alpha_{2}\mathcal{V}^{\pi}_{2}+\dots+\alpha_{d}\mathcal{V}^{\pi}_{d}$. In this line of work, there has been recent progress that achieved learning functioning policies in high-dimensional state observation MDPs (Garg et al.,, 2021). The study achieves this by learning a soft $Q$-function from observing expert demonstrations, and the study further argues that it is possible to recover rewards from the learnt soft state-action value function.